Twice as many adults as children attended a violin concert. During the intermission, 530 adults and 25 children left the concert hall. The number of children remaining in the hall became twice as many as the number of adults remaining in the hall. How many people attended the violin concert?
Answer by ikleyn(52788)

.

x children and 2x adults initially/originally.


Then, after intermission


    2*(2x - 530) = (x - 25).


Simplify and find x


    4x - 2*530 = x - 25

    4x - x = 2*530 - 25

    3x     = 1035

     x     = 1035/3 = 345.


345 children and 2*345 = 690 adults attended the concert.


ANSWER.  The total of the people attended the concert was 345 + 690 = 1035.